# Stabilized symplectic embeddings of higher-dimensional ellipsoids

### v1

@@ 4 Gravitational descendants of cotangent bundles
    The moduli space $\mathcal{M}_{\widehat{W}_{\mathrm{tor}}}^{J_{\infty}}(\gamma)$ consists of simple planes and has virtual dimension zero. Moreover, it is compact as it carries the minimal symplectic area. To explain this, note that any plane in this moduli space can be compactified to a Maslov index $2$ disk with boundary $L_{\Phi_{i},r_{i}}$ which, by the monotonicity of $L_{\Phi_{i},r_{i}}$ , must have symplectic area equal to $r_{i}/(k+1)$. A non-trivial holomorphic building that can appear as a result of degeneration in this moduli space contains two non-constant components in the top level (cf. [Faisal:2024, Theorem 6.10]), one of which can be compactified to a disk with boundary on $L_{\Phi_{i},r_{i}}$. Such a disk has a symplectic area of at least $r_{i}/(k+1)$ by the monotonicity of $L_{\Phi_{i},r_{i}}$, leaving no symplectic area for the other component. We conclude that the signed count $\mathcal{M}_{\widehat{W}_{\mathrm{tor}}}^{J_{\infty}}(\gamma)$ is well-defined and does not depend on the choice of the generic almost complex structure.
@@
    Using a suitable transversality scheme, we can assume the moduli space (5.12) is transversely cut out. For instance, following [Cieliebak2018, Section 6], one can introduce a coherent Hamiltonian perturbation supported in the interior of
@@
    Take a flat metric on $L_{\Phi,t}$. After scaling it, we can symplectically embed the codisk bundle of radius $2$, denoted by $D^{*}_{2}L_{\Phi,r}$, into $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the complement of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Perturb this metric according to [Cieliebak2018, Lemma 2.2] to a Riemannian metric $g$ such that, with respect to $g$, every closed geodesic $\gamma$ of length less than or equal to $r$ is noncontractible, nondegenerate (as a critical point of the energy functional) and satisfies
@@
    As $k\to\infty$, by SFT compactness theorem [Cieliebak_2005, MR2026549], the $J_{k}$-holomorphic sphere described in Step 01 breaks into a holomorphic building with top level in $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$, bottom level in $T^{*}L_{\Phi,t}$, and some intermediate symplectization levels in $\mathbb{R}\times S^{*}L_{\Phi,t}$.
@@
    There are at least $k+1$ $J_{\infty}$-holomorphic disks in the top level $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$, by [Faisal:2024, Lemma 6.4]. Denote these by $D_{1},D_{2},\dots,D_{k},D_{k+1}$. Since the building is the limit of the holomorphic spheres intersecting the $J_{\infty}$-holomorphic hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$ with intersection number +1+1. This means that exactly one of the components $D_{1},D_{2},\dots,D_{m}$, say $D_{m}$, of the building in the top level intersects this hypersurface; otherwise, the total intersection number will be greater than +1+1. Here, we use the fact that distinct holomorphic objects intersect positively. Therefore, the components $D_{1},D_{2},\dots,D_{m-1}$ are in the complement $B^{2k}(r)\times T^{2(n-k)}$ of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Compactifying $D_{1},D_{2},\dots,D_{m-1}$ to surfaces with boundaries on $L_{\Phi,t}$, we get
@@
    Following the arguments of [Faisal:2024, Section 6.3], we conclude that the holomorphic building has only two levels. The top level $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$ consists of $k+1$ negatively asymptotical cylindrical simple planes $D_{1},\dots,D_{k},D_{\infty}$. The bottom level $T^{*}L_{\Phi,t}$ consists of a single smooth connected punctured sphere $C_{\mathrm{bot}}$ with $k+1$ positive punctures carrying the tangency constraint at $p$. Moreover, each of the disks $D_{1},\dots,D_{k},D_{\infty}$ is of Maslov index $2$.
@@
    For each $i=1,\dots,k,\infty$, following [Faisal:2024, Section 6.6], one can glue a half-cylinder to the moduli space $\mathcal{M}^{J_{\infty}}_{[D_{i}]}(\gamma_{i})$ to create a Maslov index $2$ disk $\bar{D}_{i}:(D^{2},\partial D^{2})\to(\mathbb{CP}^{k}\times T^{2(n-k)},L_{\Phi,t})$ whose boundary passes through a fixed generic point $q\in L_{\Phi,t}$. Moreover, this disk is $J$-holomorphic for generic $\Omega$-compatible almost complex structure $J$ on $\mathbb{CP}^{k}\times T^{2(n-k)}$. Let $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{i}]}(q)$ denote the connected component of the moduli space containing $\bar{D}_{i}$. By construction, we have the non-vanishing of the signed count
@@
    This symplectic area is minimal in the sense that any non-trivial breaking in $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{\infty}]}(q)$ requires a symplectic area strictly greater than this (cf. [Faisal:2024, Theorem 6.10]). This implies the signed count above does not change if we vary $J$, $q$, or $t$ in the interval $[r,k+1]$. The same is true for other the moduli spaces $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{i}]}(q)$. The conclusion is that the disks $\bar{D}_{1},\bar{D}_{2},\dots,\bar{D}_{k},\bar{D}_{\infty}$ contribute non-trivially to the Landau–Ginzburg superpotential of $L_{\Phi,r}$ viewed as a Lagrangian torus in $(\mathbb{CP}^{k}\times\mathbb{C}^{n-k},r\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}})$. Whereas, the disks $\bar{D}_{1},\bar{D}_{2},\dots,\bar{D}_{k}$ contribute to the Landau–Ginzburg superpotential of $L_{\Phi,r}$ viewed as a Lagrangian torus in $(B^{2k}(r)\times\mathbb{C}^{n-k},\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}})$.
@@
Shah Faisal
Department of Mathematics
Uppsala University
Sweden
E-mail address: shahmath19@gmail.com, shah.faisal@math.uu.se

### v2

@@ 4 Gravitational descendants of cotangent bundles
    The moduli space $\mathcal{M}_{\widehat{W}_{\mathrm{tor}}}^{J_{\infty}}(\gamma)$ consists of simple planes and has virtual dimension zero. Moreover, it is compact as it carries the minimal symplectic area. To explain this, note that any plane in this moduli space can be compactified to a Maslov index $2$ disk with boundary $L_{\Phi_{i},r_{i}}$ which, by the monotonicity of $L_{\Phi_{i},r_{i}}$ , must have symplectic area equal to $r_{i}/(k+1)$. A non-trivial holomorphic building that can appear as a result of degeneration in this moduli space contains two non-constant components in the top level (cf. [FAI25, Theorem 6.10]), one of which can be compactified to a disk with boundary on $L_{\Phi_{i},r_{i}}$. Such a disk has a symplectic area of at least $r_{i}/(k+1)$ by the monotonicity of $L_{\Phi_{i},r_{i}}$, leaving no symplectic area for the other component. We conclude that the signed count $\mathcal{M}_{\widehat{W}_{\mathrm{tor}}}^{J_{\infty}}(\gamma)$ is well-defined and does not depend on the choice of the generic almost complex structure.
@@
    Using a suitable transversality scheme, we can assume the moduli space (5.12) is transversely cut out. For instance, following [CM18, Section 6], one can introduce a coherent Hamiltonian perturbation supported in the interior of
@@
    Take a flat metric on $L_{\Phi,t}$. After scaling it, we can symplectically embed the codisk bundle of radius $2$, denoted by $D^{*}_{2}L_{\Phi,r}$, into $\mathbb{CP}^{k}\times T^{2(n-k)}$ in the complement of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Perturb this metric according to [CM18, Lemma 2.2] to a Riemannian metric $g$ such that, with respect to $g$, every closed geodesic $\gamma$ of length less than or equal to $r$ is noncontractible, nondegenerate (as a critical point of the energy functional) and satisfies
@@
    As $k\to\infty$, by SFT compactness theorem [CM05, BEH+03], the $J_{k}$-holomorphic sphere described in Step 01 breaks into a holomorphic building with top level in $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$, bottom level in $T^{*}L_{\Phi,t}$, and some intermediate symplectization levels in $\mathbb{R}\times S^{*}L_{\Phi,t}$.
@@
    There are at least $k+1$ $J_{\infty}$-holomorphic disks in the top level $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$, by [FAI25, Lemma 6.4]. Denote these by $D_{1},D_{2},\dots,D_{k},D_{k+1}$. Since the building is the limit of the holomorphic spheres intersecting the $J_{\infty}$-holomorphic hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$ with intersection number +1+1. This means that exactly one of the components $D_{1},D_{2},\dots,D_{m}$, say $D_{m}$, of the building in the top level intersects this hypersurface; otherwise, the total intersection number will be greater than +1+1. Here, we use the fact that distinct holomorphic objects intersect positively. Therefore, the components $D_{1},D_{2},\dots,D_{m-1}$ are in the complement $B^{2k}(r)\times T^{2(n-k)}$ of the hypersurface $\mathbb{CP}^{k-1}\times T^{2(n-k)}$. Compactifying $D_{1},D_{2},\dots,D_{m-1}$ to surfaces with boundaries on $L_{\Phi,t}$, we get
@@
    Following the arguments of [FAI25, Section 6.3], we conclude that the holomorphic building has only two levels. The top level $\mathbb{CP}^{k}\times T^{2(n-k)}\setminus L_{\Phi,t}$ consists of $k+1$ negatively asymptotical cylindrical simple planes $D_{1},\dots,D_{k},D_{\infty}$. The bottom level $T^{*}L_{\Phi,t}$ consists of a single smooth connected punctured sphere $C_{\mathrm{bot}}$ with $k+1$ positive punctures carrying the tangency constraint at $p$. Moreover, each of the disks $D_{1},\dots,D_{k},D_{\infty}$ is of Maslov index $2$.
@@
    For each $i=1,\dots,k,\infty$, following [FAI25, Section 6.6], one can glue a half-cylinder to the moduli space $\mathcal{M}^{J_{\infty}}_{[D_{i}]}(\gamma_{i})$ to create a Maslov index $2$ disk $\bar{D}_{i}:(D^{2},\partial D^{2})\to(\mathbb{CP}^{k}\times T^{2(n-k)},L_{\Phi,t})$ whose boundary passes through a fixed generic point $q\in L_{\Phi,t}$. Moreover, this disk is $J$-holomorphic for generic $\Omega$-compatible almost complex structure $J$ on $\mathbb{CP}^{k}\times T^{2(n-k)}$. Let $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{i}]}(q)$ denote the connected component of the moduli space containing $\bar{D}_{i}$. By construction, we have the non-vanishing of the signed count
@@
    This symplectic area is minimal in the sense that any non-trivial breaking in $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{\infty}]}(q)$ requires a symplectic area strictly greater than this (cf. [FAI25, Theorem 6.10]). This implies the signed count above does not change if we vary $J$, $q$, or $t$ in the interval $[r,k+1]$. The same is true for other the moduli spaces $\mathcal{M}^{J}_{L_{\Phi,t},[\bar{D}_{i}]}(q)$. The conclusion is that the disks $\bar{D}_{1},\bar{D}_{2},\dots,\bar{D}_{k},\bar{D}_{\infty}$ contribute non-trivially to the Landau–Ginzburg superpotential of $L_{\Phi,r}$ viewed as a Lagrangian torus in $(\mathbb{CP}^{k}\times\mathbb{C}^{n-k},r\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}})$. Whereas, the disks $\bar{D}_{1},\bar{D}_{2},\dots,\bar{D}_{k}$ contribute to the Landau–Ginzburg superpotential of $L_{\Phi,r}$ viewed as a Lagrangian torus in $(B^{2k}(r)\times\mathbb{C}^{n-k},\omega_{\mathrm{FS}}\oplus\omega_{\mathrm{std}})$.